OpenStudy (anonymous):

How do I solve the following simultaneous equation: y = x^2 and y = 5x - 6?

OpenStudy (anonymous):

x2-5x+6=0

OpenStudy (anonymous):

y=y x2=5x-6 .. . . . x=2,3

OpenStudy (anonymous):

Can you explain it? I don't really get it

OpenStudy (anonymous):

x2-2x-3x+6=0 x(x-2)-3(x-2)=0 (x-2)(x-3)=0

OpenStudy (anonymous):

got it? eliminate y from both equations to get a quadratic equation and solve it

OpenStudy (anonymous):

Ok